# Mitigating Multimodal Hallucinations via Gradient-based Self-Reflection

### v1

@@ 6 Conclusion
In conclusion, we introduce a gradient-based self-reflection method to estimate token influence and quantitatively estimate bias severity. This estimation enables the identification of object-related visual tokens, which are then integrated into an influence-aware contrastive decoding framework. This framework effectively mitigates both text-visual and co-occurrence biases, reducing hallucinations. Our method operates without requiring additional resources such as costly fine-tuning, extra models, or data statistics. Furthermore, to reduce text-visual bias in long-generated sequences, we propose a sample-dependent stopping criterion based on the proposed visual influence.
@@ Appendix A First Order Taylor Expansion
The first order Taylor expansion of $\mathcal{F}_{\theta^{*}}$ w.r.t., $\mathbf{t}^{v},\mathbf{t}^{p}$ and, all the previous output tokens $\mathbf{y}_{i}$ with $i<m$:
$\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}\approx$ $\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v(0)},\mathbf{t}^{p(0)})_{m}+\sum^{S}_{s=1}\frac{\partial(\mathcal{F}_{\theta^{*}})_{m}}{\partial{t}^{v}_{s}}\bigg{|}_{\mathbf{t}^{v(0)}}({t}^{v}_{s}-{t}^{v(0)}_{s})$ (14)
$+\sum^{N}_{n=1}\frac{\partial(\mathcal{F}_{\theta^{*}})_{m}}{\partial{t}^{p}_{n}}\bigg{|}_{\mathbf{t}^{p(0)}}({t}^{p}_{n}-{t}^{p(0)}_{n})+\sum^{m-1}_{i=1}\frac{\partial(\mathcal{F}_{\theta^{*}})_{m}}{\partial y_{i}}\bigg{|}_{y^{(0)}_{i}}(y_{i}-y^{(0)}_{i})$
@@
where $\mathbf{g}^{v}_{ms}=\frac{\partial(\mathcal{F}_{\theta^{*}})_{m}}{\partial{t}^{v}_{s}}\Big{|}_{\mathbf{t}^{v}=\mathbf{t}^{v(0)}}$, $\mathbf{g}^{p}_{mn}=\frac{\partial(\mathcal{F}_{\theta^{*}})_{m}}{\partial{t}^{p}_{n}}\Big{|}_{\mathbf{t}^{p}=\mathbf{t}^{p(0)}}$ and $\mathbf{g}^{y}_{mi}=\frac{\partial(\mathcal{F}_{\theta^{*}})_{m}}{\partial y_{i}}\Big{|}_{\mathbf{y}_{i}=\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v(0)},\mathbf{t}^{p(0)})}$ are the first order gradients w.r.t., visual tokens $\mathbf{t}^{v}$, text prompt $\mathbf{t}^{p}$ and previous outputs $y_{i}$ and $Const$ denotes all other terms that are constant w.r.t., the $\mathbf{t}^{v},\mathbf{t}^{p}$.
@@ Appendix B Interpreting Contrastive Decoding through KL Divergence
Kullback-Leibler (KL) divergence can be used to interpret contrastive decoding, For example, in GACD VA mode, it measures the divergence between the reference distribution $p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{p},y_{<m})$ to the visual joint distribution $p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{v},\mathbf{t}^{p},y_{<m})$.
$\mathrm{D}_{KL}$ $=\sum_{c}p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{v},\mathbf{t}^{p},y_{<m})\text{log}\big{(}\frac{p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{v},\mathbf{t}^{p},y_{<m})}{p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{p},y_{<m})}\big{)}$ (15)
$=\sum_{c}p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{v},\mathbf{t}^{p},y_{<m})(\text{log}(p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{v},\mathbf{t}^{p},y_{<m}))-\text{log}(p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{p},y_{<m})))$
$=\sum_{c}p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{v},\mathbf{t}^{p},y_{<m})(\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}[c]-\text{log}(\sum\exp(\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}))$
$-\mathcal{F}_{\theta^{*}}(\mathbf{t}^{p})_{m}[c]+\text{log}(\sum\exp(\mathcal{F}_{\theta^{*}}(\mathbf{t}^{p})_{m})))$
$=\sum_{c}p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{v},\mathbf{t}^{p},y_{<m})(\underbrace{(\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}-\mathcal{F}_{\theta^{*}}(\mathbf{t}^{p})_{m})}_{\text{adjustment term}}[c]+Const),$
where $p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{v},\mathbf{t}^{p},y_{<m})=\mathbf{\sigma}(\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m})$, $p(y_{cm}|\mathcal{F}_{\theta^{\star}},\mathbf{t}^{p},y_{<m})=\mathbf{\sigma}(\mathcal{F}_{\theta^{*}}(\mathbf{t}^{p})_{m})$ and $c$ represents a class in the predefined vocabulary. The adjustment term $\mathcal{F}_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}-\mathcal{F}_{\theta^{*}}(\mathbf{t}^{p})_{m}$ increases the KL divergence, thereby emphasizing the impact of visual tokens.
@@ Appendix C Result on modern MLLMs
We further evaluate modern MLLMs, LLaVA-v1.6 and Intern-VL2, on the Amber dataset. The results in Tab. 6 demonstrate that our method consistently improves MLLM performance, even when their original performance is high.
@@ Appendix E Results on MMBench
We further evaluate our method on MMBench [33]. The results in Tab. 8 indicate that our method improves the overall performance and achieves consistent improvements across MLLMs on Coarse Perception (CP). This outcome aligns with the intended effect of our method, as its focus on increasing visual influence is directly linked to improving coarse perception capabilities. For other metrics, our method yields minor improvements due to the possible reason that certain abilities, such as Logical Reasoning (LR), rely more on the language component of MLLMs and cannot be enhanced solely by increasing visual influence.
@@ Appendix F Results on MM-Vet
The evaluation on MM-Vet [49] in Tab. 9 shows that our method achieves consistent overall (Total) improvement, along with enhancements in recognition (Rec) and Optical Character Recognition (OCR), indicating its effectiveness in improving visual recognition. However, its performance varies across other metrics, including knowledge (Know), generalization (Gen), spatial awareness (Spat), and math (Math), suggesting that our method, which focuses on token influence balancing, may not effectively enhance the generalization ability of MLLMs.
@@
Appendix G Results on MME
Our evaluation on MME [13] dataset is presented in Tab. 10. Our method achieves better overall (Total) results and equal or improved performance in existence and counting, demonstrating its effectiveness in object recognition. However, it does not improve position accuracy and exhibits varying behavior on color. This diversity may stem from the inherent capabilities of MLLMs, which cannot be solely enhanced through token influence balancing.
@@ Appendix G Results on MME
Appendix H Other Results of POPE
Table 11: More Results on POPE [29].
@@ Appendix H Other Results of POPE
Appendix I Different Sampling Strategies
Tab. 12 presents an ablation study on sampling strategies (non-greedy vs. greedy). We follow the non-greedy sampling setting of VCD [26], where both top-p and temperature are set to 1. As shown, our method consistently improves performance across both sampling strategies.
@@ Appendix I Different Sampling Strategies
Appendix J Different Model Size
@@
Appendix K Gradient Computation Details and Efficiency Ablation
Our method obtains gradients directly through PyTorch’s ‘torch.autograd.grad’ on input tokens, eliminating the need for manual derivations and facilitating straightforward reproduction. For comparison, we evaluate Integrated Gradients (IG) [11, 34, 21] using the SIG [11] implementation; Table 14 presents this ablation. In the table, “IG” denotes the SIG-based results, while “direct” refers to our torch.autograd.grad approach. Both methods yield comparable accuracy and F1-score, but the direct-gradient variant is substantially more efficient.
@@
Appendix L Norm Selection for Token Influence
@@
Appendix M Hyper parameter Study
Maximum $\alpha_{m}$. To determine the optimal maximum amplification factor ($\alpha_{m}$) and understand its impact on the model performance, we conducted a search over values from 1 to 6 using the LLaVA-v1.5 model on the POPE dataset for discriminative tasks. For open-ended generation tasks on a subset of the MSCOCO dataset following [10], we observed garbled text when $\alpha_{m}$ was set to 5; therefore, we limited our search to values from 1 to 4. As shown in Tab. 16 and Tab. 17, discriminative task on POPE is less sensitive to the value of $\alpha_{m}$. The performance on the generative task gets improved while $\alpha_{m}$ increases but later drops. Therefore, in our experiments, the optimal maximum amplification factor ($\alpha_{m}$) is set to 5 and 3 for discriminative and generation tasks, respectively.
Table 16: $\alpha_{m}$ Study For Discriminative Task On POPE [29] in MSCOCO Adversarial Setting.
@@
Early Stopping Threshold. To set the early stopping threshold properly, we conducted a search on a subset of the MSCOCO dataset subset following [10]. Recognizing that the visual influence ratio varies across models, we first analyze the sample-wise visual influence to identify an appropriate range for this study. By searching over the corresponding range, we show results in Tab. 18. These results demonstrate that varying the ES threshold primarily mediates the trade-off between recall and hallucination rate. Our goal is to have balanced recall ($R$) and instance-level hallucination ($C_{I}$), leading us to select thresholds of $7\%$ for LLaVA-v1.5 and LLaVA-v1.6, $25\%$ for InstructBLIP, $2.5\%$ for mPLUG-Owl2 and $10\%$ for InternVL2. We additionally ran an experiment to measure the ES activation rate using LLaVA-v1.5 with ES threshold $7\%$. As shown in Tab. 19, ES fires on only $8.7\%$ of the test samples, and when it does, the generated responses are on average just 0.7 tokens shorter. This indicates that ES rarely, and only minimally, truncates outputs.
@@
Appendix N Additional Implementation and Experimental Details
Noun tokens are identified using the spaCy library via its en_core_web_sm model. For experiments on the Amber dataset [44], we adopt the original data splits and evaluation metrics. In the MSCOCO [30] subset, we follow the data partitioning and evaluation protocol of Deng et al. [10], with splits available in their official repository. For the LLaVA-QA90 [32], MME [13] and POPE [29] datasets, our setup replicates that of Leng et al. [26] and use their provided scoring scripts for LLaVA-QA90. Experiments on MMBench [33], MM-Vet [49] follow the VLMEvalKit_InternVL2_5 repository. All comparison methods are executed using their official code; we only modify them to enforce greedy sampling and a uniform maximum generation length to align with our experimental settings.
Appendix O Question Category Results on the AMBER Dataset
@@
Appendix P Confidence and Visual Influence
(a) mPLUGOwl2
(b) mPLUGOwl2-GACD
Appendix Q Qualitative Example on Occluded Images
Figure 7: Example of our method applied to an occluded image.
@@
In this section, we analyze the computational cost of the proposed method. Typically, gradient computation adds an overhead equivalent to 1–2 times the cost of a forward pass. As a result, our approach increases the overall computational cost by up to five times compared to the baseline while remaining significantly lower than fine-tuning-based methods. Notably, contrastive logits generation involves only a small number of visual tokens, helping to reduce the actual runtime cost.
Appendix S Influence Ratio in Image Caption
Appendix T Influence Ratio in VQA
@@
Appendix U Limitations and Future Work
Appendix V Broader Impacts

### v2

@@ 5 Conclusion
In conclusion, we introduce a gradient-based self-reflection method to estimate token influence and quantitatively estimate bias severity. This estimation enables the identification of object-related visual tokens, which are then integrated into an influence-aware constrained decoding framework. This framework effectively mitigates both text-visual and co-occurrence biases, reducing hallucinations. Our method operates without requiring additional resources such as costly fine-tuning, extra models, or data statistics. Furthermore, to reduce text-visual bias in long-generated sequences, we propose a sample-dependent stopping criterion based on the proposed visual influence.
@@ Appendix A First Order Taylor Expansion
Let $\mathbf{z}_{m}^{\star}\in\mathbb{R}^{|\mathcal{V}|}$ denote the step-$m$ logits $\mathbf{z}_{m}^{\star}=\pi_{\theta^{\star}}(\mathbf{t}^{v},\mathbf{t}^{p},\mathbf{y}_{<m})$. Around a reference point $(\mathbf{t}^{v(0)},\mathbf{t}^{p(0)},\mathbf{y}_{<m}^{(0)})$, the detailed first-order Taylor expansion of the logit $\mathbf{z}_{m}^{\star}$ is
$\mathbf{z}_{m}^{\star}\;\approx$ $\underbrace{\mathbf{z}_{m}^{\star(0)}}_{\displaystyle\pi_{\theta^{\star}}(\mathbf{t}^{v(0)},\mathbf{t}^{p(0)},\mathbf{y}^{(0)}_{<m})}\;+\;\sum_{s=1}^{S}\,\mathbf{g}^{v}_{ms}\!\left(\mathbf{t}^{v}_{s}-\mathbf{t}^{v(0)}_{s}\right)$ (15)
$\;+\;\sum_{n=1}^{N}\,\mathbf{g}^{p}_{mn}\!\left(\mathbf{t}^{p}_{n}-\mathbf{t}^{p(0)}_{n}\right)\;+\;\sum_{i=1}^{m-1}\,\mathbf{g}^{y}_{mi}\!\left(\mathbf{y}_{i}-\mathbf{y}^{(0)}_{i}\right)$
$=$ $\sum^{S}_{s=1}\mathbf{g}^{v}_{ms}{t}^{v}_{s}+\sum^{N}_{n=1}\mathbf{g}^{p}_{mn}{t}^{p}_{n}+\sum^{m-1}_{i=1}\mathbf{g}^{y}_{mi}y_{i}$
$+\underbrace{\mathbf{z}_{m}^{\star(0)}-\sum^{S}_{s=1}\mathbf{g}^{v}_{ms}{t}^{v(0)}_{s}-\sum^{N}_{n=1}\mathbf{g}^{p}_{mn}{t}^{p(0)}_{n}-\sum^{m-1}_{i=1}\mathbf{g}^{y}_{mi}y^{(0)}_{i}}_{Const},$
@@
where the (token-wise) Jacobians are
$\mathbf{g}^{v}_{ms}\;\coloneqq\;\left.\frac{\partial\,\mathbf{z}^{\star}_{m}}{\partial\,\mathbf{t}^{v}_{s}}\right|_{\mathbf{t}^{v}=\mathbf{t}^{v(0)}},\qquad\mathbf{g}^{p}_{mn}\;\coloneqq\;\left.\frac{\partial\,\mathbf{z}^{\star}_{m}}{\partial\,\mathbf{t}^{p}_{n}}\right|_{\mathbf{t}^{p}=\mathbf{t}^{p(0)}},\qquad\mathbf{g}^{y}_{mi}\;\coloneqq\;\frac{\partial\,\mathbf{z}^{\star}_{m}}{\partial\,\mathbf{y}_{i}}\Big|_{\mathbf{y}=\mathbf{y}_{<m}^{(0)}},$ (16)
and $\mathbf{z}_{m}^{\star(0)}=\pi_{\theta^{\star}}(\mathbf{t}^{v(0)},\mathbf{t}^{p(0)},\mathbf{y}^{(0)}_{<m})$. Here $\big|_{\cdot}$ denotes evaluation at the reference point. Each $\mathbf{g}^{v}_{ms},\,\mathbf{g}^{p}_{mn},\,\mathbf{g}^{y}_{mi}$ maps a small token perturbation in its corresponding embedding space to a perturbation of the logit vector in $\mathbb{R}^{|\mathcal{V}|}$. And $Const$ denotes all other terms that are constant w.r.t., the $\mathbf{t}^{v},\mathbf{t}^{p}$.
@@ Appendix B Interpreting Contrastive Decoding through KL Divergence
Kullback-Leibler (KL) divergence can be used to interpret contrastive decoding, It measures the divergence between the reference distribution $p_{\theta^{\star}}(y_{cm}|\mathbf{t}^{o},\mathbf{t}^{p},y_{<m})$ to the $\mathbf{t}^{u}$ joint distribution $p_{\theta^{\star}}(y_{cm}|\mathbf{t}^{v},\mathbf{t}^{p},y_{<m})$, where $\mathbf{t}^{v}=\{\mathbf{t}^{u},\mathbf{t}^{o}\}$.
$\mathrm{D}_{KL}$ $=\sum_{c}p_{\theta^{\star}}(y_{cm}|\mathbf{t}^{v},\mathbf{t}^{p},y_{<m})\text{log}\big(\frac{p_{\theta^{\star}}(y_{cm}|\mathbf{t}^{v},\mathbf{t}^{p},y_{<m})}{p_{\theta^{\star}}(y_{cm}|\mathbf{t}^{o},\mathbf{t}^{p},y_{<m})}\big)$ (17)
$=\sum_{c}p_{\theta^{\star}}(y_{cm}|\mathbf{t}^{v},\mathbf{t}^{p},y_{<m})(\text{log}(p_{\theta^{\star}}(y_{cm}|\mathbf{t}^{v},\mathbf{t}^{p},y_{<m}))-\text{log}(p_{\theta^{\star}}(y_{cm}|\mathbf{t}^{o},\mathbf{t}^{p},y_{<m})))$
$=\sum_{c}p_{\theta^{\star}}(y_{cm}|\mathbf{t}^{v},\mathbf{t}^{p},y_{<m})([\pi_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}]_{c}-\text{log}(\sum\exp(\pi_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}))$
$-[\pi_{\theta^{*}}(\mathbf{t}^{o},\mathbf{t}^{p})_{m}]_{c}+\text{log}(\sum\exp(\pi_{\theta^{*}}(\mathbf{t}^{o},\mathbf{t}^{p})_{m})))$
$=\sum_{c}p_{\theta^{\star}}(y_{cm}|\mathbf{t}^{v},\mathbf{t}^{p},y_{<m})(\underbrace{[\pi_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m}-\pi_{\theta^{*}}(\mathbf{t}^{o},\mathbf{t}^{p})_{m}}_{\text{adjustment term}}]_{c}+Const),$
where $p_{\theta^{\star}}(y_{cm}|\mathbf{t}^{v},\mathbf{t}^{p},y_{<m})=\mathbf{\sigma}(\pi_{\theta^{*}}(\mathbf{t}^{v},\mathbf{t}^{p})_{m})$, $p_{\theta^{\star}}(y_{cm}|\mathbf{t}^{o},\mathbf{t}^{p},y_{<m})=\mathbf{\sigma}(\pi_{\theta^{*}}(\mathbf{t}^{o},\mathbf{t}^{p})_{m})$ and $c$ represents a class in the predefined vocabulary. The adjustment term increases the KL divergence, thereby emphasizing the impact of visual tokens.
@@ Appendix C Result on modern MLLMs
We further evaluate recent MLLMs, LLaVA-v1.6, Qwen2-VL, and Intern-VL2, on the AMBER dataset. As shown in Tab. 6, our method consistently improves performance, even when the base models already achieve strong results. Moreover, it surpasses SOTA comparison methods. This improvement stems from the model’s self-reflection ability, which effectively identifies biases in baseline MLLMs and adapts the outputs accordingly.
@@ Appendix E Results on MMBench
We further evaluate our method on MMBench Liu et al. (2025). The results in Tab. 8 indicate that our method improves the overall performance and achieves consistent improvements across MLLMs on Coarse Perception (CP). This outcome aligns with the intended effect of our method, as its focus on increasing visual influence is directly linked to improving coarse perception capabilities. For other metrics, our method yields minor improvements due to the possible reason that certain abilities, such as Logical Reasoning (LR), rely more on the language component of MLLMs and cannot be enhanced solely by increasing visual influence.
@@ Appendix F Results on MM-Vet
The evaluation on MM-Vet Yu et al. (2023) in Tab. 9 shows that our method achieves consistent overall (Total) improvement, along with enhancements in recognition (Rec) and Optical Character Recognition (OCR), indicating its effectiveness in improving visual recognition. However, its performance varies across other metrics, including knowledge (Know), generalization (Gen), spatial awareness (Spat), and math (Math), suggesting that our method, which focuses on token influence balancing, may not effectively enhance the generalization ability of MLLMs.
@@
Appendix G Results on ScienceQA and Vizwiz
We evaluate our method on two complementary multimodal benchmarks. ScienceQA Lu et al. (2022) integrates images, textual context, and curriculum knowledge, requiring models to perform structured multimodal reasoning. VizWiz Gurari et al. (2018), in contrast, consists of visual questions collected from blind users and features real-world challenges such as low-quality images, conversational queries, and unanswerable cases. These datasets jointly assess both reasoning under structured multimodal contexts and robustness in unconstrained real-world settings. As shown in Table 10, our approach consistently improves over the LLaVA-1.5 baseline. These gains demonstrate the effectiveness of our hallucination mitigation strategy in enhancing visual grounding across both knowledge-driven and real-world VQA tasks.
Table 10: Comparison of LLaVA-1.5 and our method on ScienceQA and VizWiz datasets.
MLLMs Method ScienceQA(%) $\uparrow$ VizWiz(%) $\uparrow$
LLaVA-V1.5 base 66.2 48.7
Ours 68.7 52.8
Appendix H Revisiting the Accuracy–Informativeness Trade-off
In the main paper, we report recall and output length alongside CHAIR scores, since our objective is to evaluate models under a balance of accuracy and informativeness. This choice is deliberate: our early stopping mechanism can be tuned to shorten responses, which naturally reduces CHAIR scores but at the expense of recall and content richness. Consequently, the trade-off introduced by early stopping is an explicit design choice, and it can be adjusted depending on the requirements of a specific application.
Direct comparison with SOTA methods that omit recall and generation length is therefore not entirely fair. Our analysis confirms that CHAIR scores can drop substantially when outputs are truncated, underscoring the importance of jointly reporting recall and length to present a complete view of performance. Without these complementary metrics, lower CHAIR values may simply reflect shorter, less informative responses rather than genuine improvements in visual grounding. To enable a fairer comparison with prior work, we adjust our early stopping threshold to $12\%$. Under this setting, our method achieves lower CHAIR scores while maintaining competitive recall, thereby outperforming both approaches. This demonstrates that our framework not only mitigates hallucination effectively but also preserves informativeness. Moreover, the adjustable nature of the early stopping mechanism ensures that users can flexibly select the optimal balance between accuracy and informativeness for their specific use cases.
Table 11: Comparison with SOTA Methods with $12\%$ Early Stopping Threshold.
MLLM Method $C_{s}$ $\downarrow$ $C_{i}$ $\downarrow$ R $\uparrow$ Len $\uparrow$
LLaVA-v1.5 base 48.8 13.4 78.6 99.8
PAI Liu et al. (2024c) 24.8 6.9 - -
Middle Jiang et al. (2025) 25.0 6.7 - -
Ours_ES_12% 23.5 6.5 55.0 54.1
Appendix I Results on MME
Our evaluation on MME Fu et al. (2024) dataset is presented in Tab. 12. Our method achieves better overall (Total) results and equal or improved performance in existence and counting, demonstrating its effectiveness in object recognition. However, it does not improve position accuracy and exhibits varying behavior on color. This diversity may stem from the inherent capabilities of MLLMs, which cannot be solely enhanced through token influence balancing.
@@ Appendix I Results on MME
Appendix J Other Results of POPE
Table 13: More Results on POPE Li et al. (2023b).
@@
Appendix K Different Sampling Strategies
Tab. 14 presents an ablation study on sampling strategies (non-greedy vs. greedy). We follow the non-greedy sampling setting of VCD Leng et al. (2024), where both top-p and temperature are set to 1. As shown, our method consistently improves performance across both sampling strategies.
@@
Appendix L Different Model Size
@@
Appendix M Gradient Computation Details and Efficiency Ablation
@@
Our method obtains gradients directly through PyTorch’s ‘torch.autograd.grad’ on input tokens, eliminating the need for manual derivations and facilitating straightforward reproduction. For comparison, we evaluate Integrated Gradients (IG) Enguehard (2023); Lundstrom et al. (2022); Kapishnikov et al. (2021) using the SIG Enguehard (2023) implementation; Table 16 presents this ablation. In the table, “IG” denotes the SIG-based results, while “direct” refers to our torch.autograd.grad approach. Both methods yield comparable accuracy and F1‐score, but the direct‐gradient variant is substantially more efficient.
Appendix N Norm Selection for Token Influence
@@
Appendix O Hyper parameter Study
Maximum $\alpha_{m}$. To determine the optimal maximum amplification factor ($\alpha_{m}$) and understand its impact on the model performance, we conducted a search over values from 1 to 6 using the LLaVA-v1.5 model on the POPE dataset for discriminative tasks. For open-ended generation tasks on a subset of the MSCOCO dataset following Deng et al. (2024), we observed garbled text when $\alpha_{m}$ was set to 5; therefore, we limited our search to values from 1 to 4. As shown in Tab. 18 and Tab. 19, discriminative task on POPE is less sensitive to the value of $\alpha_{m}$. The performance on the generative task gets improved while $\alpha_{m}$ increases but later drops. Therefore, in our experiments, the optimal maximum amplification factor ($\alpha_{m}$) is set to 5 and 3 for discriminative and generation tasks, respectively.
Table 18: $\alpha_{m}$ Study For Discriminative Task On POPE Li et al. (2023b) in MSCOCO Adversarial Setting.
@@
Early Stopping Threshold. To set the early stopping threshold properly, we conducted a search on a subset of the MSCOCO dataset subset following Deng et al. (2024). Recognizing that the visual influence ratio varies across models, we first analyze the sample-wise visual influence to identify an appropriate range for this study. By searching over the corresponding range, we show results in Tab. 20. These results demonstrate that varying the ES threshold primarily mediates the trade-off between recall and hallucination rate. Our goal is to have balanced recall ($R$) and instance-level hallucination ($C_{I}$), leading us to select thresholds of $7\%$ for LLaVA-v1.5 and LLaVA-v1.6, $25\%$ for InstructBLIP, $2.5\%$ for mPLUG-Owl2 and $10\%$ for InternVL2. We additionally ran an experiment to measure the ES activation rate using LLaVA-v1.5 with ES threshold $7\%$. As shown in Tab. 21, ES fires on only $8.7\%$ of the test samples, and when it does, the generated responses are on average just 0.7 tokens shorter. This indicates that ES rarely, and only minimally, truncates outputs.
@@
Appendix P Confidence and Visual Influence
(a) mPLUGOwl2
(b) mPLUGOwl2-GACD
Appendix Q Question Category Results on the AMBER Dataset
@@
In this section, we analyze the computational cost of the proposed method. On the POPE dataset, our method increases the average computational cost by $101.44\%$ in TFLOPs. This is expected for decoding-based approaches, as they require two forward passes. Importantly, the visual encoder is executed only once, and the second pass processes a limited number of input tokens. Thus, although additional computation is introduced, the overall overhead remains within a practical range.
Table 23: TFLOPs of Baseline vs. Ours with LLaVA-v1.5 on the POPE Dataset.
Method TFLOPs Relative Increase
LLaVA-v1.5 $9.68$
Ours $19.49$ $101.44\%$
Appendix S Influence Ratio in VQA
@@
Appendix T Influence Ratio in Image Caption
Appendix U Qualitative Example on Occluded Images
Figure 9: Example of our method applied to an occluded image.
Appendix V Additional Implementation and Experimental Details
Noun tokens are identified using the spaCy library via its en_core_web_sm model. For experiments on the Amber dataset Wang et al. (2023), we adopt the original data splits and evaluation metrics. In the MSCOCO Lin et al. (2014) subset, we follow the data partitioning and evaluation protocol of Deng et al. Deng et al. (2024), with splits available in their official repository. For the LLaVA-QA90 Liu et al. (2024b), MME Fu et al. (2024) and POPE Li et al. (2023b) datasets, our setup replicates that of Leng et al. Leng et al. (2024) and use their provided scoring scripts for LLaVA-QA90. Experiments on MMBench Liu et al. (2025), MM-Vet Yu et al. (2023) follow the VLMEvalKit_InternVL2_5 repository. All comparison methods are executed using their official code; we only modify them to enforce greedy sampling and a uniform maximum generation length to align with our experimental settings.
Appendix W Limitations and Future Work
Appendix X Broader Impacts